# Quantum metric and localization in a quasicrystal

### v1

@@ I Introduction
Quasicrystalline materials have since their experimental discovery in 1984 [1] been subject to growing interest thanks to their highly peculiar electronic properties. Quasicrystals are long-range ordered materials like crystals, but without translation symmetry. Instead, the long-range order in quasicrystals stems from scale-invariance of their atomic structure, visible in their diffraction pattern [1]. Consequently, the emergent electronic properties of quasicrystals are closely related to their unusual atomic structure. Recent studies have in particular focused on the localization properties of the electronic states, demonstrating their criticality [2] and topological properties [3, 4, 5, 6, 7]. There is also growing interest in how these localization properties influence interacting [8, 9] and ordered states such as superconductivity [10, 11, 12, 13, 14, 15, 16] or its proximity effect [17, 18] and Josephson effect [19], topological superconductivity [20, 21, 22, 23], and antiferromagnetism [24]. Recently, quasiperidicity has also been found in moiré structures [25, 26, 27, 28], further broadening the importance of quasicrystals.
Among the most studied quasicrystals is the Fibonacci chain [29]. It is a chain of sites (atoms) coupled by two types of hoppings that alternate following a regular, but non-crystalline pattern. Since it is one-dimensional (1D), it is simple to study and both numerical and analytical results of the electronic spectrum are available [30, 31, 32, 33]. Yet, it remains similar to more elaborate quasicrystals, as it exhibits most of the general properties of quasicrystals [34, 35, 14], including the effect of hidden dimensions [36, 37, 38]. It has therefore come to serve as an effective model to describe the essential physics of also 2D and 3D quasicrystals [29, 34, 14]. Moreover, the Fibonacci chain has been experimentally realized and studied in a wide range of physical systems [39, 40, 3, 41, 42, 43, 44, 45, 46, 35, 47].
@@
The ability of the quantum metric to measure localization and its relation to long-range phase coherence makes it an interesting tool for quasicrystals. To this end, we explore how the quantum metric can quantify localization properties of the Fibonacci chain. By showing how the quantum metric incorporates information about distances between local symmetry centers of each eigenstate and how these are inherently linked to the scale-invariance of the Fibonacci chain, we are able to establish a direct correspondence between the structural properties of the chain and the localization of its electronic states as measured by the quantum metric. The quantum metric thereby provides a well-refined measure of electronic localization, directly rooted in the scale-invariance of the chain. This makes the quantum metric a natural tool for understanding localization in general quasicrystals, also beyond the 1D Fibonacci chain, in contrast to the IPR and related tools that primarily focus only on the number of occupied sites [49, 50, 35]. Moreover, with the quantum metric being a numerically easily accessible observable, with rapidly emerging experimental signatures [60, 61, 62], it offers direct access to electronic localization of quasicrystals.
@@ II Constructing a Fibonacci chain
Approximants can alternatively be built using the ’cut-and-project’ method [63], which emphasizes the long-range structure of the chain. It consists of cutting a stripe from a tilted 2D square lattice, and projecting the sites it contains onto a 1D line, as illustrated in Fig. 1(a). When projected along the line-cut direction, the bonds along the $a$-axis result in long bonds $t_{L}$, whereas bonds along the $b$-axis project as short bonds $t_{S}$. The slope of the tilted square lattice, $\tan(\alpha)$, entirely determines the obtained chain: rational slopes correspond to periodic chains, while irrational slopes give rise to true quasicrystals. The Fibonacci chain corresponds to a slope equal to the golden mean $\tan(\alpha)=\frac{1+\sqrt{5}}{2}$. Rational slopes close to this ratio lead to the approximants $(F_{n})$.
@@ IV Quantum Metric
In Fig. 3(b), the quantum metric of each energy gap is obtained by considering all states below the gap as occupied. Despite some fluctuations, the quantum metric displays a clear structure, with both a smooth increase as a function of gap label and clear bounds. The upper bound is independent of gap label and in fact stems from the finite size of the approximant. Indeed, the spread of the states, and thus the quantum metric, can never be larger than the squared length of the chain, which thus sets the upper bound. Longer approximants would be necessary to resolve the difference between such extended bulk states and true Bloch waves with infinite quantum metric. A lower bound is also clearly present and, importantly, smoothly increases as a function of gap label, creating the overall structure. Indeed, this is not just a strict lower bound but most gaps have a quantum metric that even closely tracks this bound. This means that as the gaps get narrower (larger gap labels), the electrons consistently hybridize into states stretching over ever larger scales. However, for absolute gap labels larger than around $60$ in Fig. 3(b), the quantum metric saturates and decreases in some gaps, which we attribute to finite size effects. This is consistent with the picture of requiring larger and larger local neighborhoods to resolve states around ever narrower gaps, also resulting in edge states penetrating far into the bulk. Thus we already establish here that the quantum metric is a distinguishable, yet simply accessible, observable and probe of spatial properties of the eigenstates in quasicrystals. Moreover, our results demonstrate that, by setting the filling fraction of the electronic states, which sets the Fermi level in a specific gap, it is possible to tune the localization length of the eigenstates to range from completely trivial localization to a delocalization over the full length of the chain. Further tunability is also offered by changing the modulation strength of the chain.
@@ V Renormalization scheme and quantum metric
Thus the quantum metric measures the spread, or extent, of the dimerized state between these two groups of sites related by a mirror symmetry. As shown in Fig. 4, each gap in the spectrum opens due to the hybridization of two similar groups of sites into symmetric and antisymmetric superpositions. Then Eq. (18) gives that the quantum metric calculated in that gap includes the measure of the square distance between these two groups of sites. Thus, the smooth increase of the quantum metric as a function of gap labels shown in Fig. 3(b) is directly related to the fact that the gaps appearing deeper in the spectrum are associated to larger groups of sites, associated larger spatial separation in Eq. (18) compared to smaller (absolute valued) gap labels. This establishes the quantum metric as a natural measure of localization and spatial structure of the eigenstates of quasicrystals.
@@ VII Concluding remarks
In this work, we use the quantum metric to study the localization properties of the eigenstates in a quasicrystal, focusing on the 1D Fibonacci chain. The quantum metric probes the real-space spread of the electronic states. In contrast to the IPR, it can thus efficiently discriminate states occupying the same number of sites, but with different separation in space, while still being an easily accessible observable. As we establish, the quantum metric even captures the importance of the many local symmetry centers present within the quasicrystal. Since such local symmetry centers are concatenated throughout the chain in a scale-invariant fashion, the quantum metric is in many ways an optimal and natural probe of locality on quasicrystals. With the quantum metric being numerically easily accessible and with experimental probes of it being rapidly uncovered [62], the quantum metric thus offers great promise for understanding the physics of quasicrystals, both theoretically and experimentally.
@@
The relation between properties of a 1D quasicrystal and a 2D crystal fundamentally stems from the fact that the 1D Fibonacci chain can be built by projecting a 2D square lattice on a line, as emphasized by the cut-and-project method described in Sec. II. The cut-and-project method is in fact a general method to obtain quasicrystals, in any dimension [76, 77, 78]. Thus, the approach we develop here to understand localization can be adapted to other quasicrystals, whose spatial localization properties, expressed through the full quantum metric, are then also inherited from higher-dimensional crystals, independent of them being topological or not. This is reminiscent of the already establish relation between the properties of quasicrystals and their higher-dimensional parent crystals in terms of their topological properties [6], while here we establish it also for the localization properties.
Beyond the quest for understanding and characterizing of the peculiar localization properties of quasicrystals, our work also emphasizes the quantum metric as an indispensable tool to quantify localization in scale-invariant systems and relate it to physical observables. Indeed, the quantum metric incorporates the notion of distance in real space, which is absent in many other tools currently used to study localization, such as the IPR. Moreover, the quantum metric can be directly related to easily accessible observables, such as conductivity [79] and optical responses [80], making it an experimentally accessible probe. In the presence of electronic correlations, it has also been shown to be related to the superfluid weight in flat-bands superconductors [55, 81, 56] and the occurrence of fractional quantum Hall effect in Chern insulators [57, 58, 59]. This is an especially intriguing future direction, given the growing interest in superconducting quasicrystals [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].

### v2

@@ I Introduction
Quasicrystalline materials have since their experimental discovery in 1984 [1] been subject to growing interest thanks to their highly peculiar electronic properties. Quasicrystals are long-range ordered materials like crystals, but without translation symmetry. Instead, the long-range order in quasicrystals stems from scale-invariance of their atomic structure, visible in their diffraction pattern [1]. Consequently, the emergent electronic properties of quasicrystals are closely related to their unusual atomic structure. Recent studies have in particular focused on the localization properties of the electronic states, demonstrating their criticality [2] and topological properties [3, 4, 5, 6, 7]. There is also growing interest in how these localization properties influence interacting [8, 9] and ordered states such as superconductivity [10, 11, 12, 13, 14] or its proximity effect [15, 16] and Josephson effect [17], topological superconductivity [18, 19, 20, 21], and antiferromagnetism [22]. Recently, quasiperiodicity has also been found in moiré structures [23, 24, 25, 26], further broadening the importance of quasicrystals.
Among the most studied quasicrystals is the Fibonacci chain [27]. It is a chain of sites (atoms) coupled by two types of hoppings that alternate following a regular, but non-crystalline pattern. Since it is one-dimensional (1D), it is simple to study and both numerical and analytical results of the electronic spectrum are available [28, 29, 30, 31]. Yet, it remains similar to more elaborate quasicrystals, as it exhibits most of the general properties of quasicrystals [32, 33, 13], including the effect of hidden dimensions [34, 35, 36]. It has therefore come to serve as an effective model to describe the essential physics of also 2D and 3D quasicrystals [27, 32, 13]. Moreover, the Fibonacci chain has been experimentally realized and studied in a wide range of physical systems [37, 38, 3, 39, 40, 41, 42, 43, 44, 45, 33, 46].
@@
The ability of the quantum metric to measure localization and its relation to long-range phase coherence makes it an interesting tool for quasicrystals. To this end, we explore how the quantum metric can quantify localization properties of the Fibonacci chain. By showing how the quantum metric incorporates information about distances between local symmetry centers of each eigenstate and how these are inherently linked to the scale-invariance of the Fibonacci chain, we are able to establish a direct correspondence between the structural properties of the chain and the localization of its electronic states as measured by the quantum metric. The quantum metric thereby provides a well-refined measure of electronic localization, directly rooted in the scale-invariance of the chain. This makes the quantum metric a natural tool for understanding localization in general quasicrystals, also beyond the 1D Fibonacci chain, in contrast to the IPR and related tools that primarily focus only on the number of occupied sites [48, 49, 33]. Moreover, with the quantum metric being a numerically easily accessible observable, with rapidly emerging experimental signatures [59, 60, 61, 62], it offers direct access to electronic localization of quasicrystals.
@@ II Constructing a Fibonacci chain
Approximants can alternatively be built using the ’cut-and-project’ method [63], which emphasizes the long-range structure of the chain. It consists of cutting a stripe from a tilted 2D square lattice, and projecting the sites it contains onto a 1D line, as illustrated in Fig. 1(a). When projected along the line-cut direction, the bonds along the $a$-axis result in long bonds $t_{L}$, whereas bonds along the $b$-axis project as short bonds $t_{S}$. The slope of the tilted square lattice, $\tan(\alpha)$, entirely determines the obtained chain: rational slopes correspond to periodic chains, while quadratically irrational slopes give rise to true quasicrystals. The Fibonacci chain corresponds to a slope equal to the golden mean $\tan(\alpha)=\frac{1+\sqrt{5}}{2}$. Rational slopes close to this ratio lead to the approximants $(F_{n})$.
@@ IV Quantum Metric
In Fig. 3(b), the quantum metric of each energy gap is obtained by considering all states below the gap as occupied. Despite some fluctuations, the quantum metric displays a clear structure, with both a smooth increase as a function of gap label and clear bounds. The upper bound is independent of gap label and in fact stems from the finite size of the approximant. Indeed, the spread of the states, and thus the quantum metric, can never be larger than the squared length of the chain, which thus sets the upper bound. Longer approximants would be necessary to resolve the difference between such extended bulk states and true Bloch waves with infinite quantum metric. A lower bound is also clearly present and, importantly, smoothly increases as a function of gap label, creating the overall structure. Indeed, this is not just a strict lower bound but most gaps have a quantum metric that even closely tracks this bound. This means that as the gaps get narrower (larger gap labels), the electrons consistently hybridize into states stretching over ever larger scales. This is consistent with previous results based on Kohn’s localization length [70]. However, for absolute gap labels larger than around $60$ in Fig. 3(b), the quantum metric saturates and decreases in some gaps, which we attribute to finite size effects. This is consistent with the picture of requiring larger and larger local neighborhoods to resolve states around ever narrower gaps, also resulting in edge states penetrating far into the bulk. Thus we already establish here that the quantum metric is a distinguishable, yet simply accessible, observable and probe of spatial properties of the eigenstates in quasicrystals. Moreover, our results demonstrate that, by setting the filling fraction of the electronic states, which sets the Fermi level in a specific gap, it is possible to tune the localization length of the eigenstates to range from completely trivial localization to a delocalization over the full length of the chain. Further tunability is also offered by changing the modulation strength of the chain.
@@ V Renormalization scheme and quantum metric
Thus the quantum metric measures the spread, or extent, of the dimerized state between these two groups of sites related by a mirror symmetry. As shown in Fig. 4, each gap in the spectrum opens due to the hybridization of two similar groups of sites into symmetric and antisymmetric superpositions. Then Eq. (18) gives that the quantum metric calculated in that gap includes the measure of the square distance between these two groups of sites. Thus, the smooth increase of the quantum metric as a function of gap labels shown in Fig. 3(b) is directly related to the fact that the gaps appearing deeper in the spectrum are associated to larger groups of sites, associated larger spatial separation in Eq. (18) compared to smaller (absolute valued) gap labels. This establishes the quantum metric as a natural measure of localization and spatial structure of the eigenstates of quasicrystals. An explicit derivation of the evolution of the quantum metric under renormalization group transformations has been obtained in an independent but simultaneous work [73].
@@ VII Concluding remarks
In this work, we use the quantum metric to study the localization properties of the eigenstates in a quasicrystal, focusing on the 1D Fibonacci chain. The quantum metric probes the real-space spread of the electronic states. In contrast to the IPR, it can thus efficiently discriminate states occupying the same number of sites, but with different separation in space, while still being an easily accessible observable. As we establish, the quantum metric even captures the importance of the many local symmetry centers present within the quasicrystal. Since such local symmetry centers are concatenated throughout the chain in a scale-invariant fashion, the quantum metric is in many ways an optimal and natural probe of locality on quasicrystals. With the quantum metric being numerically easily accessible and with experimental probes of it being rapidly uncovered [62], the quantum metric thus offers great promise for understanding the physics of quasicrystals [78], both theoretically and experimentally.
@@
The relation between properties of a 1D quasicrystal and a 2D crystal fundamentally stems from the fact that the 1D Fibonacci chain can be built by projecting a 2D square lattice on a line, as emphasized by the cut-and-project method described in Sec. II. The cut-and-project method is in fact a general method to obtain quasicrystals, in any dimension [79, 80, 81]. Thus, the approach we develop here to understand localization can be adapted to other quasicrystals, whose spatial localization properties, expressed through the full quantum metric, are then also inherited from higher-dimensional crystals, independent of them being topological or not. This is reminiscent of the already establish relation between the properties of quasicrystals and their higher-dimensional parent crystals in terms of their topological properties [82, 6], while here we establish it also for the localization properties.
Beyond the quest for understanding and characterizing of the peculiar localization properties of quasicrystals, our work also emphasizes the quantum metric as an indispensable tool to quantify localization in scale-invariant systems and relate it to physical observables. Indeed, the quantum metric incorporates the notion of distance in real space, which is absent in many other tools currently used to study localization, such as the IPR. Moreover, the quantum metric can be directly related to easily accessible observables, such as conductivity [83] and optical responses [60, 84], making it an experimentally accessible probe. In the presence of electronic correlations, it has also been shown to be related to the superfluid weight in flat-bands superconductors [54, 85, 55] and the occurrence of fractional quantum Hall effect in Chern insulators [56, 57, 58]. This is an especially intriguing future direction, given the growing interest in superconducting quasicrystals [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].